OpenStudy (anonymous):

In what direction and by how many units is the graph of f(x)=5cos(4x-pi)-1 vertically and horizontally shifted?

OpenStudy (ribhu):

go down by 5 units and move left by pi.

OpenStudy (anonymous):

@satellite73

OpenStudy (ribhu):

i think i have answered @Patsfan12

OpenStudy (anonymous):

I don't think you go down 5 units

OpenStudy (anonymous):

down one on account of that \(-1\) out at the end

OpenStudy (anonymous):

as for right or left, easiest thing to do is solve \[4x-\pi=0\] for \(x\)

OpenStudy (anonymous):

you get \[x=\frac{\pi}{4}\]

OpenStudy (ribhu):

so sorry go down by 1 unit and shift right with pi units sorry @Patsfan12

OpenStudy (anonymous):

I see it now! Thanks guys!

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

Wouldn't you go up 4?

OpenStudy (anonymous):

compared to the graph of \(y=\sin(4x)\) the graph of \(y=\sin(4x-\pi)\) is shifted to the right \(\frac{\pi}{4}\) units

OpenStudy (anonymous):

here look at the picture http://www.wolframalpha.com/input/?i=5sin%284x-pi%29-1

OpenStudy (anonymous):

My answer choices are: up 4,right pi/4 Up 4, left pi/4 Down 4, left pi/4 Up 4, left pi/4

OpenStudy (anonymous):

@ribhu @satellite73

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Down 4, right pi/4?

OpenStudy (ribhu):

correct.

OpenStudy (anonymous):

Are you sure?

OpenStudy (anonymous):

Was this correct?